# Large Language Model Inference Acceleration: A Comprehensive Hardware Perspective

### v3

@@ 3.5.3. PIM/NDP
Memory Enhancement. (1) PIM-CXL. As the frequent transfers of these model parameters and intermediate values are performed over relatively slow device-to-device interconnects such as PCIe or NVLink, they become the key bottleneck for efficient acceleration of LLMs. Kim et al. (kim2024breakthrough, ) exploit PIM, which uses bank-level parallelization to provide higher internal memory bandwidth compared to traditional DRAM, resulting in a significant increase in on-DRAM compute bandwidth. In addition to achieving high capacity through Compute eXpress Link (CXL) memory expansion, CXL-PNM demonstrates performance improvements by integrating computational logic into memory products, consequently increasing internal memory bandwidth. Evaluation results show that the performance of memory-bounded LLMs was significantly improved with PIM and PNM by up to 4.5$\times$ and 4.4$\times$, respectively. CXL-PNM (park2024lpddr, ) first devises an LPDDR5X-based CXL memory architecture with 512GB of capacity and 1.1TB/s of bandwidth, which boasts 16$\times$ larger capacity and 10$\times$ higher bandwidth than GDDR6 and DDR5-based CXL memory architectures, respectively, under a module form-factor constraint. Second, it designs a CXL-PNM controller architecture integrated with an LLM inference accelerator, exploiting the unique capabilities of such CXL memory to overcome the disadvantages of competing technologies such as HBM-PIM and AxDIMM. The evaluation shows that for OPT-13B model, CXL-PNM achieves 42.68 tokens/s with 77.6W power consumption. (2) PIM-3D Stack. Sharda et al. (sharda2024accelerator, ) propose to use the capacitorless 3D stackable DRAM to store much larger LLMs compared to conventional DRAM at higher density. To reduce the intermediate data size, they propose to use a layer-wise sparsity-quantization hybrid (LSQH) algorithm, which induces sparsity based on calculations performed using low-bit quantization to reduce both the energy consumption and the data storage requirements. Finally, a 3D heterogeneously integrated accelerator is designed by stacking a 3D DRAM with logic dies designed in the 3nm technology node at 1GHz. The evaluation shows that for Llama2-13B, it achieves 163k tokens/s in prefill stage with 193W power consumption.
3.5.4. Comparison
@@
In Figure 12, for heterogeneous cooperation, power consumption ranges from 6.3W to 575W, with inference speeds between 4.9 tokens/s and 330 tokens/s. The energy efficiency ranges from 0.0145 tokens/J to 4.342 tokens/J. PowerInfer-2 (xue2024powerinfer, ) with Snapdragon 8 Gen 3 (CPU) achieves the lowest power consumption with 6.3W. SK Hynix AiMX (AiMX, ) (PIM/NDP) achieves the highest throughput with batch size 1 and the highest energy efficiency.
@@
  • For CPUs, power consumption ranges from 6.3W to 575W, with inference speeds between 4.9 tokens/s and 11.7 tokens/s, situated in the bottom part of the figure. The energy efficiency ranges from 0.0145 tokens/J to 1.857 tokens/J. In the end-side scenario, the CPU can achieve higher energy efficiency.
@@
  • For PIM/NDPs, power consumption ranges from 37W to 240W, with inference speeds between 36.34 tokens/s and 330 tokens/s, found in the upper part of the graph. The energy efficiency ranges from 0.55 tokens/J to 4.342 tokens/J, similar with edge-side CPU platforms.
3.6. Homogeneous Cooperation
3.6.1. Overview
@@
3.6.2. CPU
Distributed computing emerges as a prevalent strategy to mitigate single-node memory constraints and expedite LLM inference performance. He et al. (he2024distributed, ) propose an efficient distributed inference optimization solution for LLMs on CPUs. On four 5th Gen Intel Xeon Scalable Processors the result shows the generation speed on Qwen-72B is 7.14 tokens/s. He et al. (he2024inference, ) also propose new attention flow, SlimAttention, to reduce the KV cache size and ensure precision for efficient LLM inference on CPUs. The experimental results on Llama2-70b shows the latency of token generation is 4 tokens/s with 2 sockets and 11.4 tokens/s with 8 sockets on Intel Xeon 8563C CPUs.
3.6.3. FPGA
DFX (hong2022dfx, ) is a multi-FPGA acceleration which uses model parallelism and optimized dataflow to improve LLM inference speed in both prefilling and decoding phases. With the implementation on four Xilinx Alveo U280 FPGAs, DFX achieves about 120 tokens/s for GPT2-1.5B model.
4. Further Comparison
4.1. Setup
@@
  • For most ASIC platforms, we first define a post-silicon work (10609625, ) with an absolute inference speed as the baseline and scale the results by the peak computational performance (TOPS or GOPS).
@@
4.2. Hardware Comparison
4.2.1. Small batch size (bs=1)
In Figure 13, for CPU platforms, power consumption ranges from 3W to 575W, with inference speeds between 3 tokens/s and 50 tokens/s, located in the bottom part of the figure. The energy efficiency ranges from 0.014 tokens/J to 2.38 token/J. Additionally, we observe edge-side CPUs (including CPU SoCs) with 3W to 6W power consumption exhibit higher energy efficiency (0.544 tokens/J to 2.38 tokens/J) compared to GPUs. For GPU platforms, power consumption ranges from 40W to 700W, with inference speeds between 18 tokens/s and 194 tokens/s, situated in the middle right part of the figure. The energy efficiency ranges from 0.067 tokens/J to 0.825 token/J. For FPGA platforms, power consumption ranges from 45W to 225W, with inference speeds between 40 tokens/s and 333 tokens/s, also in the middle part of the figure. The energy efficiency ranges from over 0.178 tokens/J to 1.85 tokens/J, which is higher than GPU platforms and similar with edge-side CPUs. For ASIC platforms, power consumption ranges from 6.3W to 600W, with inference speeds between 9.173 tokens/s and 1800 tokens/s, found in the upper left part of the figure. The energy efficiency ranges from 0.108 tokens/J to over 18.557 tokens/J, outperforming CPU, GPU and FPGA platforms. For PIM/NDP platforms, power consumption ranges from 11.516W to 240W, with inference speeds between 42.68 tokens/s and 1998 tokens/s, located in the upper left part of the figure. The energy efficiency ranges from 0.53 tokens/J to 46.66 tokens/J, outperforming other hardware platforms.
Figure 13. LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with different optimization methods.
Figure 14. LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 8) vs power on different platforms with different optimization methods.
4.2.2. Large batch size (bs=8)
@@
4.3. Optimization Method Comparison
4.3.1. Small batch size (bs=1)
We then compare different optimization methods on the same platforms. In Figure 13, the methods employed on CPU platforms include quantization, sparsity, and heterogeneous cooperation. Quantization can achieve higher absolute inference speeds, reaching approximately 50 tokens/s, while sparsity and heterogeneous cooperation achieve speeds of 16.3 tokens/s and 11.7 tokens/s, respectively. Quantization and heterogeneous cooperation can achieve higher energy efficiency with 2.38 tokens/J and 1.86 tokens/J, respectively, compared to 0.16 tokens/J for sparsity. On GPU platforms, the methods used include quantization, sparsity, fast decoding, and operator optimization. Quantization, fast decoding, and operator optimization can achieve higher absolute inference speeds, reaching approximately 194 tokens/s, 169.68 tokens/s, and 145.04 tokens/s, respectively, while sparsity achieves only about 50 tokens/s. Regarding efficiency, these three methods show relatively small differences in performance compared to sparsity (0.16 tokens/J), with efficiencies of 0.825 tokens/J, 0.587 tokens/J, and 0.46 tokens/J, respectively. On FPGA platforms, the methods employed include quantization and sparsity, which are often used together. The highest on-board deployed speed reaches 92.5 tokens/s, with an efficiency of up to 1.32 tokens/J. Some works estimate that FPGA can achieve a maximum inference speed of 333 tokens/s and an efficiency of 1.85 tokens/J. On ASIC platforms, the methods employed include quantization, sparsity, and operator optimization. Operator optimization can achieve the highest absolute inference speeds, reaching approximately 1800 tokens/s. In comparison, sparsity and quantization achieve speeds of 42.664 tokens/s and 161.086 tokens/s, respectively. In terms of efficiency, quantization and operator optimization offer higher energy efficiency with 7.434 tokens/J and 18.557 tokens/J, respectively. Sparsity shows lower energy efficiency with only 1.421 tokens/J. On PIM/NDP platforms, the main methods used are quantization, fast decoding, and heterogeneous cooperation. Quantization can achieve higher throughput with 1998 tokens/s while fast decoding and heterogeneous cooperation only achieve 174.018 tokens/s and 330 tokens/s, respectively. In terms of efficiency, quantization can achieve higher energy efficiency with 46.66 tokens/J while fast decoding and heterogeneous cooperation only achieve 9.94 tokens/J and 4.34 tokens/s, respectively.
4.3.2. Large batch size (bs=8)
@@
5. Further Discussion
@@
Figure 15. The trends of edge AI system: multimodality, inference-time computation, and higher inference energy efficiency.
5.1. Multimodality
@@
As shown in Fig. 16, we summarize the release dates and the number of training text tokens for typical LLMs over the past five years. The number of training tokens has been growing at a rate of 2.5$\times$ per year. GPT-3 is trained using only 0.4 trillion text tokens, while Qwen2.5 and Llama3.1 use 18 trillion and 15 trillion tokens, respectively. According to the result of Epoch AI, the total number of human-generated public text tokens is $\sim$300 trillion (villalobos2024run, ). At this growth rate, it is expected that by 2027, we will face a shortage of available public text tokens for training LLMs. In addition to text tokens, there is a vast amount of other modalities of tokens on the Internet, such as image and video tokens (total $\sim$1650 trillion (villalobos2024run, )). Therefore, from the perspective of token availability, by leveraging multiple modalities of tokens for training LLMs, we can continue the LLMs’ scaling law (kaplan2020scaling, ). From the perspective of human interaction, humans rely on various senses as gateways to interact with the physical world, which is inherently multimodal. Recent research has also proven that multimodal learning provably performs better than single modal due to the more accurate estimate of the latent space representation (huang2021makes, ). These factors indicate that future LLMs should have the capability of multimodal interactions to perceive, process, and generate information across multiple modalities.
Figure 16. The number of training text tokens for typical single modal LLMs in the past 5 years. We will lack the public text tokens after 2027. By using multimodal tokens for training, we can continue the LLMs’ scaling law.
@@
Currently, more than 80% multimodal LLMs focus on both vision and language tasks (zhang2024mm, ). As shown in Fig. 17, their development can be roughly divided into three stages. First, from May to September 2023, the model size is small and the accuracy ranges from 20% to 40%. Then, from December 2023 to April 2024, with the release of GPT4v, the model accuracy significantly improves, reaching nearly 60%. After May 2024 with the release of GPT4o, more multimodal models are released and the accuracy reaches almost 80%. It should be noted that the development of multimodal LLMs (MLLMs) heavily relies on the construction of multimodal datasets. And another significant trend in MLLMs is the concept of native multimodality. By aligning diverse modalities during the training stage, native MLLMs enable end-to-end inputs and outputs.
@@
Figure 17. The timeline of multimodal (Vision-Language) LLMs: average accuracy, publication date, and model size. The results are collected from OpenCompass (OpenCompass, ).
5.2. Inference-time Compute
@@
In the LLM era, data is the fossil fuel, while the age of pre-training will end because the data is not growing (llyanips, ). Inference-time compute, as represented by OpenAI’s o1 paradigm (openai_o1, ), is now seen as the critical pathway to enhance model capabilities in the next LLM era. The principles of o1 can be illustrated through its distinction from GPT4, as shown in Fig. 18(a). During the training stage, GPT4 follows the conventional workflow: pre-training a base model followed by fine-tuning through reinforcement learning from human feedback (RLHF). In contrast, o1 introduces several innovations in the post-training stage. Beyond incorporating reinforcement learning (RL), o1 places particular emphasis on safety optimization to ensure that o1 is equipped for more nuanced tasks and safe decision-making during deployment. The difference between GPT4 and o1 becomes more pronounced in the inference stage. GPT4 generates answers directly without explicitly involving iterative reasoning. By comparison, o1 introduces a “Think&Summary” phase before producing answers. This step leverages advanced techniques such as Chain-of-Thought (CoT) reasoning and tree search to perform more intricate logical reasoning and step-by-step problem solving. Compared to GPT4o, the o1 paradigm achieves a 50% improvement in accuracy on the complex logical reasoning tasks like AI Math Evaluation (AIME) test dataset. However, the inference cost is increased by 10-100$\times$ (o1inferencecost, ).
@@
Inference-time compute can enhance the model capabilities but also substantially increase inference time, particularly for complex reasoning tasks. As shown in Fig. 18(b), in the GSM8K benchmark, we employ Llama2-7B as the backbone LLM, combined with the open-source LLM-Reasoners (hao2024llm, ). The inference time is increased by 678$\times$ compared to using Llama2-7B directly. This substantial time overhead primarily stems from the iterative multi-step process inherent in inference-time compute, which is necessary to satisfy the requirements of more complex reasoning task. Therefore, choosing the appropriate scenarios and conditions for employing inference-time compute remains a topic worthy of further research. Additionally, inference-time compute also changes the runtime breakdown of the LLM inference. As shown in Fig. 18(c), the proportion of the prefill stage significantly increases from 1.5% to 23.5%, while the decode stage decreases from 98.5% to 54.8%. The increase in the prefill stage is partly due to the need to merge the human input tokens with the template input tokens as the whole LLM input, which effectively extends the input sequence length. Moreover, in the iterative process, intermediate output tokens are cumulatively fed back as inputs, leading to a gradual increase in the sequence length for prefill. Meanwhile, the introduction of the Process Reward Model (PRM) accounts for 21.7% of the time. This shift of runtime breakdown highlights that future optimization methods for inference-time compute should consider all prefill stage, decode stage, and PRM simultaneously.
@@
Figure 18. Inference-time Compute: (a) the difference of training and inference stages, (b) the increment of inference time, and (c) runtime breakdown of inference time.
5.3. Higher Inference Energy Efficiency
In future edge applications, particularly in high real-time scenarios such as embodied intelligence represented by robotics control and autonomous driving, the system’s control frequency needs to reach 100-1000 Hz (stewart1990implementing, ; bena2023hybrid, ; mafi2009parallel, ; atkins2023mintnet, ). This implies that systems using LLMs as the decision-making center for actions or instructions should achieve an inference speed of at least 100-1000 tokens/s. Since edge devices need to consider energy supply, the power of edge chips should be <20W (the power consumption of human brain (markram2012human, ; magistretti2015cellular, )) and then the inference efficiency should exceed 10 tokens/J. However, as shown in the Fig. 19, current commercial edge chips such as Tesla RPU (teslarpu, ) and NVIDIA Jetson Orin (nvidiaorin, ), even with software optimization methods, only achieve an actual inference speed of 10-30 tokens/s and an efficiency of <1 tokens/J. This presents a gap of 1-2 orders of magnitude in both absolute inference speed and efficiency compared to future edge application demands.
To achieve this goal, we should continue the algorithm-hardware co-design approach. On the algorithm side, more aggressive model compression methods can be employed with high-quality datasets for retraining to minimize accuracy loss. On the hardware side, in addition to designing specific hardware units to support algorithm optimizations, it is necessary to design more innovative chip architectures, such as 3D DRAM-stacked integrated architectures that significantly reduce data transfer latency and energy consumption, dataflow architectures that optimize data transfer paths to improve computational efficiency, and wafer-scale architectures that provide higher computational and storage capabilities through ultra-large-scale integration. These co-design approaches can effectively bridge the current gaps in inference speed and energy efficiency of edge-side chips, meeting the demands of future applications.
Figure 19. The performance gap between the future and the current edge AI systems.
6. Conclusion

### v4

@@ 3.5.5 PIM/NDP
Memory Enhancement. (1) PIM-CXL. As the frequent transfers of these model parameters and intermediate values are performed over relatively slow device-to-device interconnects such as PCIe or NVLink, they become the key bottleneck for efficient acceleration of LLMs. Kim et al. [273] exploit PIM, which uses bank-level parallelization to provide higher internal memory bandwidth compared to traditional DRAM, resulting in a significant increase in on-DRAM compute bandwidth. In addition to achieving high capacity through Compute eXpress Link (CXL) memory expansion, CXL-PNM demonstrates performance improvements by integrating computational logic into memory products, consequently increasing internal memory bandwidth. Evaluation results show that the performance of memory-bounded LLMs was significantly improved with PIM and PNM by up to 4.5$\times$ and 4.4$\times$, respectively. CXL-PNM [275] first devises an LPDDR5X-based CXL memory architecture with 512GB of capacity and 1.1TB/s of bandwidth, which boasts 16$\times$ larger capacity and 10$\times$ higher bandwidth than GDDR6 and DDR5-based CXL memory architectures, respectively, under a module form-factor constraint. Second, it designs a CXL-PNM controller architecture integrated with an LLM inference accelerator, exploiting the unique capabilities of such CXL memory to overcome the disadvantages of competing technologies such as HBM-PIM and AxDIMM. The evaluation shows that for OPT-13B model, CXL-PNM achieves 42.68 tokens/s with 77.6W power consumption. (2) PIM-3D Stack. Sharda et al. [165] propose to use the capacitorless 3D stackable DRAM to store much larger LLMs compared to conventional DRAM at higher density. To reduce the intermediate data size, they propose to use a layer-wise sparsity-quantization hybrid (LSQH) algorithm, which induces sparsity based on calculations performed using low-bit quantization to reduce both the energy consumption and the data storage requirements. Finally, a 3D heterogeneously integrated accelerator is designed by stacking a 3D DRAM with logic dies designed in the 3nm technology node at 1GHz. The evaluation shows that for Llama2-13B, it achieves 163k tokens/s in prefill stage with 193W power consumption. PIM-AI [281] stacks computational units on DDR5/LPDDR5 memory chips to improve performance and energy efficiency. Simulations show that PIM-AI achieves 30 tokens/s and 35.97 tokens/J for Llama2-7B. Lue et al. [170] propose a Flash-based Computing-in-Memory architecture using 3D NOR for accelerating LLM inference. For Llama-7B with W4A8, they can achieve $>$200 tokens/s inference speed.
3.5.6 Comparison
@@
In Figure 12, for heterogeneous cooperation, power consumption ranges from 6.3W to 880W, with inference speeds between 4.9 tokens/s and 1998 tokens/s. The energy efficiency ranges from 0.0057 tokens/J to 46.66 tokens/J. PowerInfer-2 [260] with Snapdragon 8 Gen 3 (CPU) achieves the lowest power consumption with 6.3W. Guo et al. [163] (PIM/NDP) achieves the highest throughput with batch size 1 and the highest energy efficiency.
@@
  • For CPUs, power consumption ranges from 6.3W to 880W, with inference speeds between 4.9 tokens/s and 22.73 tokens/s, situated in the bottom part of the figure. The energy efficiency ranges from 0.0057 tokens/J to 2.381 tokens/J. In the end-side scenario, the CPU can achieve higher energy efficiency.
  • For GPUs, TightLLM is the only one using heterogeneous cooperation, which is similar to CPU platforms.
  • For FPGAs, GLITCHES is the only one using heterogeneous cooperation, which achieves higher throughput than CPU/GPU platforms.
@@
  • For PIM/NDPs, power consumption ranges from 11.52W to 300W, with inference speeds between 36.34 tokens/s and 1998 tokens/s, found in the upper part of the graph. The energy efficiency ranges from 0.787 tokens/J to 46.66 tokens/J, outperforming other platforms.
3.6 Homogeneous Cooperation
3.6.1 Overview
@@
3.6.2 CPU
Distributed computing emerges as a prevalent strategy to mitigate single-node memory constraints and expedite LLM inference performance. He et al. [284] propose an efficient distributed inference optimization solution for LLMs on CPUs. On four 5th Gen Intel Xeon Scalable Processors the result shows the generation speed on Qwen-72B is 7.14 tokens/s. He et al. [285] also propose new attention flow, SlimAttention, to reduce the KV cache size and ensure precision for efficient LLM inference on CPUs. The experimental results on Llama2-70b shows the latency of token generation is 4 tokens/s with 2 sockets and 11.4 tokens/s with 8 sockets on Intel Xeon 8563C CPUs.
3.6.3 FPGA
DFX [286] is a multi-FPGA acceleration which uses model parallelism and optimized dataflow to improve LLM inference speed in both prefilling and decoding phases. With the implementation on four Xilinx Alveo U280 FPGAs, DFX achieves about 120 tokens/s for GPT2-1.5B model. In this paper, we propose LoopLynx, a scalable dataflow architecture for efficient LLM inference that optimizes FPGA usage through a hybrid spatial-temporal design. The design of LoopLynx incorporates a hybrid temporal-spatial architecture, where computationally intensive operators are implemented as large dataflow kernels. This achieves high throughput similar to spatial architecture, and organizing and reusing these kernels in a temporal way together enhances FPGA peak performance. Furthermore, to overcome the resource limitations of a single device, we provide a multi-FPGA distributed architecture that overlaps and hides all data transfers so that the distributed accelerators are fully utilized. By doing so, LoopLynx can be effectively scaled to multiple devices to further explore model parallelism for large-scale LLM inference. Evaluation of GPT2 model demonstrates that LoopLynx can achieve comparable performance to state-of-the-art single FPGA-based accelerations. In addition, compared to Nvidia A100, our accelerator with a dual-FPGA configuration delivers a 2.52x speed-up in inference latency while consuming only 48.1% of the energy.
3.6.4 ASIC
Chen et al. [287] proposes an agile framework to support two homogeneous binary and floating-point PE arrays. Experimental results illustrate it achieves $\sim$80 tokens/s for Llama-6.7B with 5.72W.
4 Further Comparison
4.1 Setup
@@
  • For most ASIC platforms, we first define a post-silicon work [150] with an absolute inference speed as the baseline and scale the results by the peak computational performance (TOPS or GOPS). We use the open-source cycle-based accurate hardware simulator for BitMoD [153] to simulate the performance of the baselines like OliVe [148].
@@
Figure 13: LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with different optimization methods.
4.2 Hardware Comparison
4.2.1 Small batch size (bs=1)
In Figure 13, for CPU platforms, power consumption ranges from 3W to 880W, with inference speeds between 3 tokens/s and 50 tokens/s, located in the bottom part of the figure. The energy efficiency ranges from 0.0057 tokens/J to 2.38 token/J. Additionally, we observe edge-side CPUs (including CPU SoCs) with 3W to 6W power consumption exhibit higher energy efficiency (0.544 tokens/J to 2.38 tokens/J) compared to GPUs. For GPU platforms, power consumption ranges from 40W to 700W, with inference speeds between 18 tokens/s and 343.4 tokens/s, situated in the middle right part of the figure. The energy efficiency ranges from 0.06 tokens/J to 0.8586 token/J. For FPGA platforms, power consumption ranges from 2.49W to 572W, with inference speeds between 3.61 tokens/s and 450 tokens/s, also in the middle part of the figure. The energy efficiency ranges from over 0.1537 tokens/J to 6.304 tokens/J, which is higher than GPU platforms and similar with edge-side CPUs. For ASIC platforms, power consumption ranges from 1.632W to 600W, with inference speeds between 9.95 tokens/s and 2700 tokens/s, found in the upper left part of the figure. The energy efficiency ranges from 0.1075 tokens/J to over 13.99 tokens/J, outperforming CPU, GPU and FPGA platforms. For PIM/NDP platforms, power consumption ranges from 11.516W to 300W, with inference speeds between 10 tokens/s and 1998 tokens/s, located in the upper left part of the figure. The energy efficiency ranges from 0.1429 tokens/J to 46.66 tokens/J, outperforming other hardware platforms.
Figure 14: LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 8) vs power on different platforms with different optimization methods.
4.2.2 Large batch size (bs=8)
@@
4.3 Optimization Method Comparison
4.3.1 Small batch size (bs=1)
We then compare different optimization methods on the same platforms. In Figure 13, the methods employed on CPU platforms include quantization, sparsity, and heterogeneous cooperation. Quantization can achieve higher absolute inference speeds, reaching approximately 50 tokens/s, while sparsity and heterogeneous cooperation achieve speeds of 16.3 tokens/s and 22.73 tokens/s, respectively. Quantization and heterogeneous cooperation can achieve higher energy efficiency with 2.38 tokens/J, compared to 0.16 tokens/J for sparsity. On GPU platforms, the methods used include quantization, sparsity, fast decoding, and operator optimization. Quantization, fast decoding can achieve higher absolute inference speeds, reaching approximately 343.4 tokens/s and 274.4 tokens/s, respectively, while operator optimization and sparsity achieve only about 181 tokens/s and 76.69 tokens/s. Regarding efficiency, these three methods show relatively small differences in performance compared to sparsity (0.434 tokens/J), with efficiencies of 0.859 tokens/J, 0.686 tokens/J, and 0.522 tokens/J, respectively. On FPGA platforms, the methods employed include quantization, sparsity and heterogeneous cooperation, which are often used together. The highest on-board deployed speed reaches 290 tokens/s, with an efficiency of up to 6.304 tokens/J. Some works achieve a performance of 450 tokens/s with an energy efficiency of 0.87 tokens/J through heterogeneous GPUs and FPGAs. On ASIC platforms, the methods employed include quantization, and operator optimization. Operator optimization can achieve the highest absolute inference speeds, reaching approximately 2700 tokens/s. In comparison, quantization achieves speeds of 161.1 tokens/s. In terms of efficiency, quantization and operator optimization offer similar and high energy efficiency with 8.49 tokens/J and 12.34 tokens/J, respectively. On PIM/NDP platforms, the main methods used are quantization, fast decoding, and heterogeneous cooperation. Quantization can achieve higher throughput with 1998 tokens/s while fast decoding and heterogeneous cooperation only achieve 174.018 tokens/s and 400 tokens/s, respectively. In terms of efficiency, quantization also achieves higher energy efficiency with 46.66 tokens/J while fast decoding and heterogeneous cooperation only achieve 9.94 tokens/J and 4.34 tokens/s, respectively.
4.3.2 Large batch size (bs=8)
@@
5 Further Discussion
Figure 15: The trends of edge AI system: multimodality, inference-time computation, and higher inference energy efficiency.
@@ 5 Further Discussion
5.1 Multimodality
@@
As shown in Fig. 16 left, we summarize the release dates and the number of training text tokens for typical LLMs over the past five years. The number of training tokens has been growing at a rate of 2.5$\times$ per year. GPT-3 is trained using only 0.4 trillion text tokens, while Qwen2.5 and Llama3.1 use 18 trillion and 15 trillion tokens, respectively. According to the result of Epoch AI, the total number of human-generated public text tokens is $\sim$300 trillion [305]. At this growth rate, it is expected that by 2027, we will face a shortage of available public text tokens for training LLMs. In addition to text tokens, there is a vast amount of other modalities of tokens on the Internet, such as image and video tokens (total $\sim$1650 trillion [305]). Therefore, from the perspective of token availability, by leveraging multiple modalities of tokens for training LLMs, we can continue the LLMs’ scaling law [60]. From the perspective of human interaction, humans rely on various senses as gateways to interact with the physical world, which is inherently multimodal. Recent research has also proven that multimodal learning provably performs better than single modal due to the more accurate estimate of the latent space representation [306]. These factors indicate that future LLMs should have the capability of multimodal interactions to perceive, process, and generate information across multiple modalities.
@@
Currently, more than 80% multimodal LLMs focus on both vision and language tasks [307]. As shown in Fig. 16, their development can be roughly divided into three stages. First, from May to September 2023, the model size is small and the accuracy ranges from 20% to 40%. Then, from December 2023 to April 2024, with the release of GPT4v, the model accuracy significantly improves, reaching nearly 60%. After May 2024 with the release of GPT4o, more multimodal models are released and the accuracy reaches almost 80%. It should be noted that the development of multimodal LLMs (MLLMs) heavily relies on the construction of multimodal datasets. And another significant trend in MLLMs is the concept of native multimodality. By aligning diverse modalities during the training stage, native MLLMs enable end-to-end inputs and outputs.
@@
Figure 16: The timeline of multimodal (Vision-Language) LLMs: average accuracy, publication date, and model size. The results are collected from OpenCompass [308].
5.2 Inference-time Compute
@@
In the LLM era, data is the fossil fuel, while the age of pre-training will end because the data is not growing [309]. Inference-time compute, as represented by OpenAI’s o1 paradigm [310], is now seen as the critical pathway to enhance model capabilities in the next LLM era. The principles of o1 can be illustrated through its distinction from GPT4, as shown in Fig. 17(a). During the training stage, GPT4 follows the conventional workflow: pre-training a base model followed by fine-tuning through reinforcement learning from human feedback (RLHF). In contrast, o1 introduces several innovations in the post-training stage. Beyond incorporating reinforcement learning (RL), o1 places particular emphasis on safety optimization to ensure that o1 is equipped for more nuanced tasks and safe decision-making during deployment. The difference between GPT4 and o1 becomes more pronounced in the inference stage. GPT4 generates answers directly without explicitly involving iterative reasoning. By comparison, o1 introduces a “Think&Summary” phase before producing answers. This step leverages advanced techniques such as Chain-of-Thought (CoT) reasoning and tree search to perform more intricate logical reasoning and step-by-step problem solving. Compared to GPT4o, the o1 paradigm achieves a 50% improvement in accuracy on the complex logical reasoning tasks like AI Math Evaluation (AIME) test dataset. However, the inference cost is increased by 10-100$\times$ [311].
@@
Inference-time compute can enhance the model capabilities but also substantially increase inference time, particularly for complex reasoning tasks. As shown in Fig. 17(b), in the GSM8K benchmark, we employ Llama2-7B as the backbone LLM, combined with the open-source LLM-Reasoners [312]. The inference time is increased by 678$\times$ compared to using Llama2-7B directly. This substantial time overhead primarily stems from the iterative multi-step process inherent in inference-time compute, which is necessary to satisfy the requirements of more complex reasoning task. Therefore, choosing the appropriate scenarios and conditions for employing inference-time compute remains a topic worthy of further research. Additionally, inference-time compute also changes the runtime breakdown of the LLM inference. As shown in Fig. 17(c), the proportion of the prefill stage significantly increases from 1.5% to 23.5%, while the decode stage decreases from 98.5% to 54.8%. The increase in the prefill stage is partly due to the need to merge the human input tokens with the template input tokens as the whole LLM input, which effectively extends the input sequence length. Moreover, in the iterative process, intermediate output tokens are cumulatively fed back as inputs, leading to a gradual increase in the sequence length for prefill. Meanwhile, the introduction of the Process Reward Model (PRM) accounts for 21.7% of the time. This shift of runtime breakdown highlights that future optimization methods for inference-time compute should consider all prefill stage, decode stage, and PRM simultaneously.
@@
Figure 17: Inference-time Compute: (a) the difference of training and inference stages, (b) the increment of inference time, and (c) runtime breakdown of inference time.
5.3 Higher Inference Energy Efficiency
In future edge applications, particularly in high real-time scenarios such as embodied intelligence represented by robotics control and autonomous driving, the system’s control frequency needs to reach 100-1000 Hz [313, 314, 315, 316]. This implies that systems using LLMs as the decision-making center for actions or instructions should achieve an inference speed of at least 100-1000 tokens/s. Since edge devices need to consider energy supply, the power of edge chips should be <20W (the power consumption of human brain [317, 318]) and then the inference efficiency should exceed 10 tokens/J. However, as shown in the Fig. 18, current commercial edge chips such as Tesla RPU [319] and NVIDIA Jetson Orin [320], even with software optimization methods, only achieve an actual inference speed of 10-30 tokens/s and an efficiency of <1 tokens/J. This presents a gap of 1-2 orders of magnitude in both absolute inference speed and efficiency compared to future edge application demands.
To achieve this goal, we should continue the algorithm-hardware co-design approach. On the algorithm side, more aggressive model compression methods can be employed with high-quality datasets for retraining to minimize accuracy loss. On the hardware side, in addition to designing specific hardware units to support algorithm optimizations, it is necessary to design more innovative chip architectures, such as 3D DRAM-stacked integrated architectures [321] that significantly reduce data transfer latency and energy consumption, dataflow architectures that optimize data transfer paths to improve computational efficiency, and wafer-scale architectures that provide higher computational and storage capabilities through ultra-large-scale integration. These co-design approaches can effectively bridge the current gaps in inference speed and energy efficiency of edge-side chips, meeting the demands of future applications.
Figure 18: The performance gap between the future and the current edge AI systems.
6 Conclusion